OpenStudy (anonymous):

Secant RM intersects secant RN at point R. Find the length of RQ .If necessary, round to the hundredths place.

Directrix (directrix):

Attached is the theorem for this problem.

Directrix (directrix):

RP * RN = RQ * RM 4 * 9 = x * ( x + 9) where x = RQ

Directrix (directrix):

Solve this for x and you should have the value of RQ 4 * 9 = x * ( x + 9)

Directrix (directrix):

@ang2016